# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ Abstract
Context. Photometric variability is a defining characteristic of young stellar objects (YSO), which can be traced back to a range of physical processes taking place at different stages of young stars’ formation and early evolution. The Gaia third Data Release (GDR3) has provided an unprecedented dataset of photometric time series, including 79 375 light curves for sources classified as YSO candidates. Thanks to its all-sky coverage, Gaia provides a unique opportunity for large-scale studies of YSO variability.
Aims. Our goal was to characterise the GDR3 sample of YSO variables to better identify the recurrence of YSO variability modes (due to accretion, extinction, rotation modulation, etc.). We make a pilot study of the applicability of the asymmetry (M) and periodicity (Q) variability metrics to characterise YSO variability with Gaia light curves. By adapting the Q&M-metrics for sparse and long-term light curves, we seek to bridge the gap between low and high-cadence survey insights on YSO variability.
Methods. We adapted the Q&M methodology for Gaia. Through refined sample selection, we identified sources with appropriate sampling for the Q&M methodology. We used the Generalised Lomb Scargle periodogram and structure functions to derive characteristic timescales of variability.
Results. We successfully derived Q&M indices for 23 000 sources in the GDR3 YSO sample. These variables were then classified into eight variability morphological classes. We linked morphological classes with physical mechanisms by using H as a proxy of accretion and -indices to gauge circumstellar material’s presence.
Conclusions. We demonstrate that the Q&M metrics can be successfully applied to study Gaia’s sparse time-series. We applied it successfully to distinguish between the various variability modes of YSOs. While our results are generally consistent with previous high-cadence, short-term studies, we find that GDR3’s long timespan yields an enhanced variety of variability mechanisms.
@@
Stars: variables: general – Stars: variables: T Tauri, Herbig Ae/Be – Accretion, accretion disks – Stars: formation – Stars: pre-main sequence Stars: rotation1 Introduction
Gaia (Gaia Collaboration et al., 2016, 2023b) is a cornerstone mission by ESA, in operation since mid-2014. In 2022, the Gaia Data Release 3 (GDR3) provided photometric and astrometric observations for 1.8 billion stars. GDR3 included epoch photometry covering 34 months of the mission provided for 11 754 237 sources, among which were variable (stellar and extragalactic Eyer et al., 2023) and all sources in the field of view (FOV) of the Gaia Andromeda Photometric Survey (GAPS Evans et al., 2023).
@@
The GDR3 variables were classified by the General Variability Detection (GVD) path into 25 classes using a supervised machine learning classifier (Rimoldini et al., 2023) with 79,375 of them classified as Young Stellar Objects (YSOs). This sample of YSO candidates was validated by Marton et al. (2023) by cross-matching with existing YSO catalogues from optical and infrared surveys, revealing that 40,000 sources were newly identified YSOs. In this paper, we further explore the GDR3 YSO variable sample by providing a more detailed look at their variability modes.
@@
Variability is a prevalent characteristic of YSOs, observed throughout their entire electromagnetic spectrum (Joy, 1945; Herbst et al., 1994; Stelzer et al., 2007; Carpenter et al., 2001; Morales-Calderón et al., 2011; Venuti et al., 2015; Roquette et al., 2020). This variability can be traced back to a wide range of physical mechanisms that take place in the formation and evolution of young stars, including mass accretion, star-disk interaction, disk dissipation, and outflows, among others (e.g. Fischer et al., 2023). Understanding YSO variability is thus fundamental to advancing our knowledge of how stars form and evolve during their first million years.
@@
Attempts to devise variability classification schemes capable of grasping the assortment of observed YSO light curves have been made since the 1990s (e.g. Herbst et al., 1994). Although light-curve characteristics such as amplitude, timescale, and periodicity were recurrent ingredients, these earlier classification schemes were limited to the specific data for which they were designed. Based on an initial sample of 162 YSOs in the young cluster NGC 2264, Cody et al. (2014) proposed a new morphology classification scheme for YSO light curves. This scheme stems from two variability metrics designed to quantify the level of asymmetry (M-index) and periodicity (Q-index) in the light curves and allow for sorting sources into seven variability classes: strictly periodic, eclipsing binary, quasi-periodic symmetric, quasi-periodic dippers, aperiodic dipper, burster, and stochastic; with two more classes, multi-periodic, and long-trend added to include sources with characteristics missed by the default Q&M classification. Previous research comparing Q&M classification results with proxies of young stellar evolution (e.g., disk IR excess and accretion-induced emission lines such as H) helped to associate these classes with various physical mechanisms behind the variability of YSOs with an emerging scenario where strictly periodic light curves are attributed to magnetically induced cold spots on the stellar surface viewed from high-inclination angles (e.g. Cody et al., 2022). Higher-amplitude quasi-periodic symmetric light curves are associated with accretion-induced hot spots at the stellar surface (e.g. Venuti et al., 2015; Cody & Hillenbrand, 2018). Dipper variability is linked to the AA Tauri phenomenon (e.g. Bouvier et al., 1999; Alencar, S. H. P. et al., 2010), where magnetic star-disk-interaction is thought to lift dusty disk material above the disk’s midplane, leading to periodic or aperiodic occultation. Bursting variables have been correlated with observables indicating intense ongoing accretion, suggesting a common origin (e.g. Stauffer et al., 2014). Stochastic variability is thought to originate from accretion related processes combined with low viewing angles (Cody & Hillenbrand, 2018). While the Q&M classification scheme was initially designed aiming applications to high-cadence space-based light curves from the Convection, Rotation and planetary Transits satellite (CoRoT Auvergne et al., 2009) and Spitzer telescope, it has since been employed in the context of varied epoch surveys including other high-cadence surveys such as Kepler mission K2 (Cody & Hillenbrand, 2018; Venuti et al., 2021; Cody et al., 2022), but also ground-based sources such as ZTF (Zwicky Transient Facility; Hillenbrand et al., 2022; Wang et al., 2023; Jiang & Hillenbrand, 2024), ASAS-SN (Bredall et al., 2020), and VVVX (Ordenes-Huanca et al., 2024).
@@
In this paper, we present the results of a pilot study that addressed the applicability of the Q&M metrics to GDR3 light curves. Applying this methodology is challenging due to the sparse and unevenly spaced observational window resulting from Gaia’s scanning law. Most sources have only 30 to 100 epoch observations spanning 900 days, which requires a complete adaptation of the original methodology introduced by Cody et al. (2014), based on a high-cadence 40-day-long light curve with more than 4,000 epochs. Still, Gaia offers all-sky coverage in contrast with the modest sample of a few hundred YSOs studied by high-cadence surveys, which also only focused on a handful of close-by star-forming regions. Hence, beyond the caveats of its sparse sampling, Gaia provides us the opportunity to derive general statistics of YSO variability from a more extensive and diverse sample. Additionally, the longer time coverage of Gaia allows us to probe longer timescales of variability, which are relevant for our understanding of the scope and prevalence of YSO variability mechanisms.
@@
This paper is organised as follows. In Sect. 2 we describe the data used in this paper. Our first goal is to test the applicability of Q&M indices to GDR3 data. For that, in Sect. 3 we perform several tests to remove sources or epoch observations that might be affected by instrumental biases. We filtered the YSO sample by removing sources with too few data points, which might have spurious variability induced by the scan angle. Finally, we split the sample in two, depending on the variability level. The high-variability sample was used to benchmark our methodology. In Sect. 4, we describe our process to implement Q&M with GDR3, with timescales derived using generalised Lomb-Scargle periodograms or structure functions. We also discuss the cases where no reliable timescale is found or those where the timescale is too long compared to the full light-curve duration. Sect. 5 presents the results on the Q&M-index, timescales, and morphological classes. We put the Q&M-index in relation with disk properties and the evolutionary stage of our samples, using the index and H emission line. In Sect. 6, we discuss our results and present a few light curves from each class, discussing the typical timescale, amplitude, and physical mechanisms behind them. We further discuss the limitations of our methods and how our results compare to previous high-cadence and ZTF studies. Finally, we summarise our work in Sect. 7.
@@ 2.1 Epoch data
- YSO sample:
- GAPS:
The Gaia Andromeda Photometric Survey (GAPS Evans et al., 2023) released epoch photometry for over 1.2 million sources located in a region of 5.5 radius centred on Andromeda. GAPS included the release of epoch photometry for both variable and sources not identified as variables in DR3, showcasing the full power of Gaia’s epoch data. The existence of stars with no variability detected by Gaia in GAPS allowed us to benchmark for our variability investigations.
Due to Gaia’s scanning law, the number of transits per source highly depends on the sky position. This is illustrated in Fig. 1. For the YSO sample, the number of epochs ranged from 14 to 138, and from 5 to 110 for GAPS. On average, YSO light curves have more epoch observations than GAPS sources.
The YSO sample and the source properties (mean magnitude, RA, DEC, etc.) were retrieved from the Gaia ESA Archive . Light curves were retrieved using the astroquery.gaia
package (Ginsburg et al., 2019). For this study, we focused our analysis on -band epoch data, as our primary interest lies in classifying variability types rather than YSO types. While using data from other bands (, ) might help distinguish between specific YSO processes, it falls beyond the scope of this work.
@@ 2.2 Supplementary data
In Sect. 3.3, we use data from GDR3 spurious variability table (vari_spurious_signals
) that was described in Holl et al. (2023) to flag sources affected by scan-angle spurious variability.
2.2.1 Epoch data from previous literature
In Sect. 6, we compare our results with previous literature investigating YSO variability using K2 and ZTF light curves. In Cody et al. (2022), K2 light curves were employed to study sources in the Taurus star-forming region, with 61 sources in common with our YSO sample; their light curves were provided to us through private communication with the authors. Hillenbrand et al. (2022), Wang et al. (2023) and Jiang & Hillenbrand (2024) used ZTF light curves to study YSO variability in the North America and Pelican Nebulae Complex, the Perseus Molecular Cloud, and the Mon R2 region, respectively. These studies included 509 sources in our YSO sample. ZTF light curves for these sources were downloaded from the ZTF archive (Masci et al., 2019) through the ZTF data access tool at the NASA/IPAC Infrared Science Archive. The total timespan of these light curves was limited to match the data released used in these previous studies, and the pre-processing of light curves followed the steps described in the previous studies.
@@ 2.2.2 YSO properties
To further validate our results in Sect. 6, we also matched our list of sources with catalogues with measurements related to the youth and evolutionary stages of YSOs. For measurements related to the line, a proxy of accretion, we used:
NEMESIS (New Evolutionary Model for Early Stages of Stars with Intelligent Systems) Catalogue of YSOs for the Orion Star Formation Complex (Roquette et al., 2025) provides a curation of parameters related to the youth of sources in that region, providing measurements for 5 231 sources in the YSO sample.
GALAH DR4 (The Galactic Archaeology with HERMES Survey Data Release 4; Buder et al., 2024) includes EW() measured from high-resolution spectroscopy for 2 085 sources in our catalogue.
LAMOST DR10 v2.0 (The General Survey of the Large Sky Area Multi-Object Spectroscopic Telescope Data; Zhao et al., 2012) includes EW() measured from low-resolution spectra for 2 172 sources.
Gaia DR3 includes spectroscopic data products based on low-resolution spectra from the Blue and Red Photometers (De Angeli et al., 2023), which provides measurements of EW() with pseudo-equivalent widths estimated as part of the Extended Stellar Parametriser for Emission-Line Stars (ESP-ELS Creevey et al., 2023) for 56 964 sources in the YSO sample.
@@ 3 Data pre-processing
In this Section, we discuss the pre-processing and filtering steps to ensure only the most reliable light curves were kept in our sample. Fig. 2 summarises our filtering process workflow. Our first cleaning step was to remove epoch observations identified by GDR3 variability processing (Eyer et al., 2023) and flagged as rejected_by_variability.
@@ 3.1 Contamination by extragalactic sources
In GAPS, we excluded from our analysis sources flagged with a probability over 0.8 of being a galaxy or a quasar, based on the Gaia DR3 classification probabilities 111from GDR3 table gaia_source classprob_dsc_combmod_galaxy and classprob_dsc_combmod_quasar (Gaia Collaboration et al., 2023a). In contrast, only six sources were similarly removed from the YSO sample.
@@ 3.2 Minimum number of epochs
We performed tests (see App. A.1) to determine the minimum number of epochs in the light curve required to ensure statistically robust variability indices. All tests were carried out as a function of magnitude to account for the different levels of photometric noise in the sample. Based on these tests, we rejected light curves with fewer than 62 observations for G average magnitudes between 9.5 and 13.5 and 37 for those outside this range. This filtering step left 44 918 sources in our YSO sample (57% of the full catalogue). In GAPS, this step left 524 096 sources (42%).
@@ 3.3 Spurious variability
Holl et al. (2023) has identified spurious variability due to scan-angle effects (combination of Gaia’s mirror aspect ratio, spin-axis precession, and orbital period, Gaia Collaboration et al., 2016). Attempts to model and filter out the spurious frequencies caused by these scan-angle effects would require assumptions such as the underlying variability signal having a sinusoidal waveform, which are not reliable priors for YSO variability. Hence, instead of attempting to mask these signals, we have followed Gaia’s data processing literature (Holl et al., 2023; Distefano et al., 2022) to identify and remove sources likely affected by spurious scan-angle signals (see App. A.2). Identifying and removing sources likely affected by spurious scan-angle signals allowed us to select 34,793 YSOs as reliable light curves (44% of the full catalogue). These sources constitute the full sample for this analysis, with some removals discussed in Sect. 4.2 due to residual overdensities caused by spurious scan angle effects identified after computing the timescales. In GAPS, our final sample contains 479 936 sources after this step (39%).
3.4 Outliers removal
Filtering large outliers is an important pre-processing step for studying the degree of asymmetry in the light curves as outliers can bias the M-index calculation (see Sect. 4.1). When dealing with high-cadence data, Cody et al. (2014) addressed this issue by 5-clipping the light curves. In the case of CoRoT and K2, as the light curves were densely sampled, this approach was unlikely to remove relevant variability features. However, due to the sparsity of the GDR3 light curves, this type of filtering needs to be addressed more carefully to avoid removing relevant variability features from the light curve. We have done so in App. A.3, finding that 636 sources in our sample had at least one outlier removed. After this outlier removal, we also verified that all affected sources still fulfilled the minimum number of epochs required for light curve analysis.

### v2

# Exploring the variability of young stars with Gaia DR3 light curves

@@ Abstract
Context. Photometric variability is a defining characteristic of young stellar objects (YSOs) that can be traced back to a range of physical processes that occur at different stages in the formation and early evolution of young stars. The Gaia third Data Release (GDR3) has provided an unprecedented dataset of photometric time series, including 79 375 light curves for sources classified as YSO candidates. Through its all-sky coverage, Gaia provides a unique opportunity for large-scale studies of YSO variability.
Aims. Our goal was to characterise the GDR3 sample of YSO variables to better identify the recurrence of YSO variability modes (caused by accretion, extinction, rotation modulation, etc.). We made a pilot study of the applicability of the asymmetry (M) and periodicity (Q) variability metrics to characterise YSO variability with Gaia light curves. By adapting the Q-M metrics for sparse and long-term light curves, we sought to bridge the gap between low- and high-cadence survey insights on YSO variability.
Methods. We adapted the Q-M method for Gaia. Through a refined sample selection, we identified sources with an appropriate sampling for the Q-M method. We used the generalised Lomb Scargle periodogram and structure functions to derive characteristic variability timescales.
Results. We successfully derived Q-M indices for 23 000 sources in the GDR3 YSO sample. These variables were then classified into eight variability morphological classes. We linked the morphological classes with physical mechanisms using H as a proxy of accretion and indices to gauge whether circumstellar material was present.
Conclusions. We demonstrate that the Q-M metrics can be successfully applied to study the sparse time series of Gaia. We applied it successfully to distinguish between the various variability modes of YSOs. While our results are generally consistent with previous high-cadence short-term studies, the long GDR3 time span yields a larger variety of variability mechanisms.
@@
Stars: variables: general – Stars: variables: T Tauri, Herbig Ae/Be – Accretion, accretion discs – Stars: formation – Stars: pre-main sequence Stars: rotation1 Introduction
@@
Gaia (Gaia2016; Gaia2023j) is a cornerstone mission by ESA that is in operation since mid-2014. In 2022, the Gaia Data Release 3 (GDR3) provided photometric and astrometric observations for 1.8 billion stars. GDR3 included epoch photometry covering 34 months of the mission provided for 11 754 237 variable sources (stellar and extragalactic; Eyer2023) and all sources in the field of view (FOV) of the Gaia Andromeda Photometric Survey (GAPS Evans23). The variables were classified into 25 classes with the GDR3 general variability detection path via supervised machine learning (Rimoldini2023), and 79 375 of them were classified as young stellar objects (YSOs). This sample of YSO candidates was validated by Marton2023 by cross-matching it with existing YSO catalogues from optical and infrared surveys. This revealed that 40 000 sources were newly identified YSOs. In this paper, we further explore the GDR3 YSO variable sample by studying their variability modes in more detail.
@@
Variability is a prevalent characteristic of YSOs, which are observed throughout their entire electromagnetic spectrum (Joy1945; Herbst94; Stelzer2007A&A...468..463S; Carpenter2001; MoralesCalderon2011ApJ...733...50M; Venuti2015A&A...581A..66V; Roquette2020A&A...640A.128R). This variability can be traced back to a wide range of physical mechanisms that occur in the formation and evolution of young stars, including mass accretion, star-disc interaction, disc dissipation, and outflows (e.g. Fischer23). Understanding YSO variability is thus fundamental to advancing our knowledge of how stars form and evolve during their first million years.
@@
Attempts to devise variability classification schemes capable of grasping the assortment of observed YSO light curves have been made since the 1990s (e.g. Herbst94). Although light curve characteristics such as amplitude, timescale, and periodicity were recurrent ingredients, these earlier classification schemes were limited to the specific data for which they were designed. Based on an initial sample of 162 YSOs in the young cluster NGC 2264, Cody_14 proposed a new morphology classification scheme for YSO light curves. This scheme stems from two variability metrics designed to quantify the level of asymmetry (M index) and periodicity (Q index) in the light curves and allow for sorting sources into seven variability classes: strictly periodic (P), eclipsing binary (EB), quasi-periodic symmetric (QPS), quasi-periodic dippers (QPD), aperiodic dipper (APD), burster (B), and stochastic (S). Two more classes, multi-periodic and long trend (L), were added to include sources with characteristics that were missed by the default Q-M classification. Previous research comparing Q-M classification results with proxies of young stellar evolution (e.g. disc IR excess and accretion-induced H emission) helped to associate these classes with various physical mechanisms behind the variability of YSOs. A scenario emerged according to which P light curves are attributed to magnetically induced cold spots on the stellar surface viewed from high-inclination angles (e.g. Cody_2022). Higher-amplitude QPS light curves are associated with accretion-induced hot spots at the stellar surface (e.g. Venuti2015A&A...581A..66V; Cody_2018). Dipper variability is linked to the AA Tauri phenomenon (e.g. Bouvier1999A&A...349..619B; Alencar2010), where magnetic star-disc-interaction is thought to lift dusty disc material above the disc mid-plane, which leads to periodic or aperiodic occultation. Bursting variables have been correlated with observation of intense ongoing accretion, suggesting a common origin (e.g. Stauffer2014AJ....147...83S). S variability is thought to originate from accretion-related processes combined with low viewing angles (Cody_2018). While the Q-M classification scheme was initially designed with the aim to apply it to high-cadence space-based light curves from the Convection, Rotation and planetary Transits satellite (CoRoT CoRoT2009A&A...506..411A) and Spitzer telescope, it has since been employed in the context of varied epoch surveys including other high-cadence surveys such as the Kepler mission K2 (Cody_2018; Venuti_2021; Cody_2022), but also ground-based surveys such as the Zwicky Transient Facility (ZTF; Hillenbrand2022; Wang2023; JiangHillenbrand2024), the All-Sky Automated Survey for Supernovae (ASAS-SN; Bredall2020), and the Vista Variables in the Vía Láctea Extended (VVVX; 2024MNRAS.533..841O).
@@
We present the results of a pilot study that addressed the applicability of the Q-M metrics to GDR3 light curves. It is challenging to apply this method because the observational window resulting from the Gaia scanning law is sparse and unevenly spaced. Most sources have only 30 to 100 epoch observations that span 900 days, which requires a full adaptation of the original method by Cody_14, based on 40-day-long high-cadence light curves with more than 4 000 epochs. Still, Gaia offers all-sky coverage, which is in contrast to the modest sample of a few hundred YSOs that were studied by high-cadence surveys focused on a handful of nearby star-forming regions. Beyond the caveats of its sparse sampling, Gaia therefore provides the opportunity to derive general statistics of YSO variability from a more extensive and diverse sample. Additionally, the longer time coverage of Gaia allowed us to probe longer variability timescales, which are relevant for our understanding of the scope and prevalence of YSO variability mechanisms.
@@
This paper is organised as follows. Sect. 2 describes the data we used. With the goal of testing whether the Q-M indices are applicable to GDR3 data, we performed several tests that we describe in Sect. 3 to identify and discard sources or epoch observations that might be affected by instrumental biases. The clean sample was divided into two variability levels, and the highest-variability sample was used to benchmark our method. In Sect. 4 we describe our implementation of the Q-M indices for GDR3, with timescales derived using generalised Lomb-Scargle periodograms or structure functions. We also discuss the cases without reliable timescale or those where the timescale is too long compared to the full light curve duration. Sect. 5 presents the results for the Q-M indices, timescales, and morphological classes. We also look into the Q-M indices in relation to the disc properties and evolutionary stage of our samples, and we use the index and the H emission line. In Sect. 6 we discuss our results and present example light curves from each class, and we discuss the typical timescale, amplitude, and physical mechanisms behind them. We further discuss the limitations of our methods and compare our results to previous high-cadence and ground-based studies. Finally, we summarise our findings in Sect. 7.
@@ 2.1 Epoch data
YSO sample: composed of 79 375 variable sources identified as YSO candidates as part of the GDR3 variability processing (Eyer2023; Rimoldini2023; Marton2023).
GAPS: The Gaia Andromeda Photometric Survey (GAPS Evans23) released epoch photometry for over 1.2 million sources located in a region of 5.5 radius centred on Andromeda. GAPS included the release of epoch photometry for all sources, i.e. both variables and non-variables, showcasing the full power of the Gaia epoch data.
Because of the Gaia scanning law, the number of transits per source highly depends on the sky position. This is illustrated in Fig. 1. For the YSO sample, the number of epochs ranged from 14 to 138, and from 5 to 110 for GAPS. On average, YSO light curves have more epoch observations than GAPS sources. General source properties (mean magnitude, coordinates, etc.) were retrieved from the GaiaArchive. Light curves were retrieved using the astroquery.gaia package (AstroqueryGinsburg2019AJ....157...98G). For this study, we focus our analysis on the -band epoch data with the goal of classifying variability types. Analysis of other bands (, ) may help distinguish between specific YSO processes, but are beyond our current scope.
@@ 2.2 Supplementary data
In Sect. 3.3, we use data from GDR3 spurious variability table (vari_spurious_signals; Holl2023) to identify sources affected by scan-angle spurious variability. This data was retrieved directly from the GaiaArchive.
2.2.1 Epoch data from the literature
In Sect. 6, we compare our results with previous literature investigating YSO variability using K2 and ZTF light curves. In Cody_2022, K2 light curves were employed to study sources in the Taurus star-forming region, with 61 sources in common with our YSO sample; light curves were provided to us through private communication with the authors. Hillenbrand2022, Wang2023, and JiangHillenbrand2024 used ZTF light curves to study YSO variability in the North America and Pelican Nebulae Complex, the Perseus Molecular Cloud, and the Mon R2 region, respectively. These studies included 509 sources in our YSO sample. ZTF light curves for these sources were downloaded from the ZTF archive (Masci2019PASP..131a8003M) through the ZTF data access tool at the NASA/IPAC Infrared Science Archive. The time span of these light curves was trimmed to match the data release used in these previous studies, and the pre-processing of light curves followed the steps described in these studies.
@@ 2.2.2 YSO properties
To support validation of our results in Sect. 6, we matched our sources with catalogues that provide measurements related to the youth and evolutionary stages of YSOs. For measurements related to the line, a proxy of accretion, we used:
The NEMESIS (New Evolutionary Model for Early Stages of Stars with Intelligent Systems) Catalogue of YSOs for the Orion star formation complex (roquette2024) compiles youth indicators for sources in that region, including measurements for 5,231 objects in our YSO sample.
GALAH DR4 (Galactic Archaeology with HERMES Survey Data Release 4; Buder2024arXiv240919858B) includes EW() from high-resolution spectroscopy for 2 085 of our sources.
LAMOST DR10 v2.0 (The General Survey of the Large Sky Area Multi-Object Spectroscopic Telescope Data; Zhao2012RAA....12..723Z) includes EW() measured from low-resolution spectra for 2 172 sources.
The Gaia-ESO Survey (Gilmore2022AA...666A.120G; Randich2022AA...666A.121R) DR5 5.1 catalogue (Hourihane2023AA...676A.129H) provides EW() and full widths at 10% of the peak flux measurements from high-resolution spectra for 751 and 765 sources, respectively.
Gaia DR3 includes spectroscopic data products based on low-resolution spectra from the blue and red photometers (BPRPspectra2023AA...674A...2D), providing measurements of EW() with pseudo-equivalent widths estimated as part of the Extended Stellar Parametriser for Emission-Line Stars (ESP-ELS Apsis2023AA...674A..26C) for 56 964 sources.
@@ 3 Data pre-processing
Fig. 2 summarises the pre-processing and filtering steps we employed to ensure that only the most reliable light curves were kept in our sample. Our first cleaning step was to remove likely spurious epoch observations based on GDR3 flag rejected_by_variability, which provides information on transits considered as outliers or extreme values and rejected as part of GDR3 variability processing steps (see sect. 3.1 of Eyer2023).
@@ 3.1 Contamination by extragalactic sources
We excluded sources from the GAPS sample which were flagged with a probability over 80% of being a galaxy or a quasar, based on the Gaia DR3 classification probabilities classprob_dsc_combmod_galaxy and classprob_dsc_combmod_quasar (GDR3_Extragalactic_2023A&A...674A..41G; Delchambre2023A&A...674A..31D). Six sources were similarly removed from the YSO sample.
@@ 3.2 Minimum number of epochs
In App. A.1 we performed tests to determine the minimum number of epochs required in the light curve to ensure statistically robust variability indices. All tests were carried out as a function of magnitude to account for the different levels of photometric noise in the sample. Based on these tests, we rejected light curves with fewer than 62 observations for mag and 37 for outside this range. This filtering step yield 44 918 sources in our YSO sample (57% of the full catalogue) and 524 096 sources in the GAPS sample (42%).
@@ 3.3 Spurious variability
Holl2023 investigated spurious variability in GDR3 light curves due to scan-angle effects (combination of the Gaia mirror aspect ratio, spin-axis precession, and orbital period, Gaia2016). Attempts to model and filter out the spurious frequencies caused by these scan-angle effects would require assumptions such as the underlying variability signal having a sinusoidal waveform, which are not reliable priors for YSO variability. Hence, instead of attempting to mask these signals, we followed the Gaia data-processing literature (Holl2023; Distefano2022) to identify and remove sources likely affected by spurious scan-angle signals (see App. A.2). After this step, we deemed 34 793 light curves in the YSO sample as reliable for further analysis (44% of the full catalogue), and 479 936 in the GAPS sample (39%).
3.4 Outlier removal
Filtering large outliers is an important for studying the degree of asymmetry in the light curves, as outliers can bias the M index calculation (see Sect. 4.1). When dealing with high-cadence data, Cody_14 addressed this issue by 5-clipping the light curves. With the densely sampled light curves of CoRoT and K2, this approach was unlikely to remove relevant variability features. However, due to the sparsity of the GDR3 light curves, this type of filtering needs to be addressed more carefully. We have done so in App. A.4, where 636 sources in our sample had at least one outlier removed. All affected sources still fulfilled the minimum number of epochs requirement after outlier removal.